OpenStudy (anonymous):

A lump of chromium (Cr) has a mass of 92.5 grams and a temperature of 89.5oC. It is placed into a calorimeter with 75.2 mL of water at 20.5oC. After stirring, the final temperature of the water, Cr metal, and calorimeter is 27.4oC. What is the specific heat of Cr metal? How do you write this out into equation?

OpenStudy (anonymous):

i still dont understand -_-

OpenStudy (anonymous):

@callisto @experimentX @hero i think u all can help?

OpenStudy (experimentx):

specific heat is defined as the amount of heat required to change the temperature by 1 degrees/

OpenStudy (experimentx):

ah ... i fotgot to add, to an unit mass of that element.

OpenStudy (experimentx):

sine you know the difference in temperature of water, volume of water(which means mass too), you can calculate the amount of heat gained by water. (as you know the sp heat cap of water is well known) => the same heat is lost by Cr => you know the mass => the amount of heat lost => the drop in temp => find sp heat const.

OpenStudy (callisto):

The basic formula used here is E=mcΔT Energy gained by water = mc(of water) ΔT = (75.2/1000) (4200)(27.4-20.5) = 2179.296J Energy loss by Cr = Energy gained by water m(Cr) x c(Cr) x ΔT = 2179.296 (92.5/1000) xc(Cr) x(89.5-27.4) = 2179.296 c(Cr) = 2179.296/5.74425 =379.4 Jkg^-1oC^-1